OpenStudy (anonymous):

Standard form equation for the parabola with the focus: F(0,4) and directrix y=3

OpenStudy (anonymous):

a parabola is a trajectory such that the from every point on the parabola the distance to the focus and normal distance to the directrix is same

OpenStudy (anonymous):

so here the focus is (0,4) and the directrix is y=3.. @Andy2Guns, are you here?

OpenStudy (anonymous):

anyway, from any arbitrary point (x,y), what is the normal distance to the line y=3? obviously |y-3| what is the distance of the point (0,4) from (x,y)? obviously \[\sqrt{(x-0)^2+(y-4)^2}\] so \[\sqrt{(x-0)^2+(y-4)^2}\]=|y-3| square in both sides and solve

OpenStudy (anonymous):

this is the equation of your required parabola